Seafoam:

Pyramid A is a square pyramid with a base side length of 12 inches and a height of 8 inches. Pyramid B has a volume of 3,456 cubic inches. How many times bigger is the volume of pyramid B than pyramid A? Give your answer as a percentage. Provide an explanation and proof for your answer to receive full credit.

simsharrison:

in my head i feel like i know how to do this but I dont want to put it out there just incase im wrong yk sooooo @AZ

Seafoam:

I think he left it on, so it says he's online. He hasn't responded to messages or anything

simsharrison:

oop- @snowflake0531

Seafoam:

Oh lord I'm done for

AZ:

\(\text{Volume of square pyramid} = \dfrac{1}{3}s^2 h\) where s is the side and h is the height so our s = 12 and h = 8 what is the volume of pyramid A?

Seafoam:

Is it 384?

AZ:

Yup! So how many times is pyramid B bigger than pyramid A? Just have to divide 3456/ 384 =

Seafoam:

9

AZ:

yeah and they want it as a percent so multiply by 100%

Seafoam:

900%..?

AZ:

yup!

Seafoam:

What about Provide an explanation and proof for your answer to receive full credit.

AZ:

show your work using that formula The volume of Pyramid A is 384 show that Pyramid B volume / Pyramid A volume * 100 = 900% and so the volume of Pyramid B is 9 00% times larger than the volume of Pyramid A.

Seafoam:

Is that all it needs?

AZ:

It's math so like I doubt they can expect any different kind of explanation- as long as you show your work

Seafoam:

What about the proof part?

AZ:

the proof is the work that you show you plugging in the numbers into the formula

Seafoam:

900%  The volume of Pyramid A is 384 Pyramid B volume / Pyramid A volume * 100 = 900% and so the volume of Pyramid B is 9 00% times larger than the volume of Pyramid A. Is that good?

AZ:

well include the formula V = 1/3 s^2 * h s = 12 h = 8 and you can even show the numbers you get in between before you reach the final answer

Seafoam:

Ok how about this 900%  V = 1/3 s^2 * h s = 12 h = 8 V= 1 / 3 12^2 * 8 The volume of Pyramid A is 384 Pyramid B volume / Pyramid A volume * 100 = 900% and so the volume of Pyramid B is 900% times larger than the volume of Pyramid A.

AZ:

might want to put a little * between 1/3 and 12^2 so you can tell that it's multiplication so V = 1/3 s^2 * h s = 12 h = 8 V= 1/3 * 12^2 * 8 The volume of Pyramid A is 384 Pyramid B volume / Pyramid A volume * 100% 3456 / 384 * 100% = 900% The volume of Pyramid B is 900% times larger than the volume of Pyramid A.

Seafoam:

Ok thank you!

AZ:

No problem!